# Conformal Prediction Assessment: A Framework for Conditional Coverage Evaluation and Selection

### 28 Mar 2026 v1

@@ 1.2 Our Contributions
    A rigorous, flexible and informative framework for evaluating conditional coverage. The reliability estimator $\hat{\eta}(x)$ can leverage any powerful method of estimating the conditional probability function and thus overcome the limitations of traditional binning methods, allowing for scalable and objective assessment of conditional validity in high-dimensional feature spaces without requiring manual stratification. Beyond a single summary statistic, we decompose the conditional coverage assessment into Safety (undercoverage risk) and Efficiency (overcoverage cost), offering nuanced insights into model failure modes. We further introduce the Conditional Validity Profile (CVP) curve, a visualization tool that characterizes the full distribution of local conditional coverage probabilities.
@@
    Reliability-driven model selection (CC-Select). Departing from the standard focus on average interval length, we propose CC-Select, a robust algorithm for selecting models based on their conditional validity. We demonstrate through extensive experiments that CC-Select consistently identifies models with superior conditional coverage performance. In addition, the reliability estimators obtained during selection are retained and used to estimate conditional coverage at new inputs.
@@ 1.3 Related Work
Conditional Coverage in Conformal Prediction.
Recent work has explored multiple approaches to improving or approximating conditional coverage in conformal prediction. One prominent line of work, initiated by Romano et al. (2019) with Conformalized Quantile Regression (CQR), directly models the conditional quantiles to produce adaptive intervals effective against heteroscedasticity. A second, broader strategy leverages an estimate of the full conditional distribution $P_{Y|X}$. This includes early work based on density estimates (Lei et al., 2013), as well as more flexible modern approaches that reframe the problem via discretization of the output space (Sesia and Romano, 2021) or by using localized approaches that up-weight calibration points near the test point (Guan, 2023). Recognizing the theoretical limitations of localized methods, Hore and Barber (2025) proposed RLCP, which achieves exact marginal coverage and, in a randomized sense, provides conditional coverage guarantees. Most recently, Gibbs et al. (2025) bridged the gap between marginal and conditional validity via a framework that interpolates between the two. Specifically, they employ augmented quantile regression on conformity scores to yield finite-sample coverage guarantees over a user-specified class of covariate shifts or subgroups. While existing methods offer theoretical improvements, their empirical validation is often limited to coarse, low-dimensional binning tests. Such approaches are inherently unscalable and may fail to detect subtle multivariate deviations. This highlights a crucial methodological gap: the lack of a scalable, quantitative, and objective framework for evaluating and comparing the conditional coverage properties of different methods. Our work is designed to directly address this issue.
Model Selection in Conformal Prediction.
Model selection in conformal prediction has begun to receive systematic attention only recently. Yang and Kuchibhotla (2025) explored the strategies of selection and aggregation and were the first to propose valid algorithmic procedures that provably maintain marginal coverage. Liang et al. (2024) built on Yang and Kuchibhotla (2025) by introducing ModSel-CP and ModSel-CP-LOO methods that correct for model selection bias in conformal prediction without requiring additional data splitting, ensuring finite-sample validity through full conformal frameworks. While these methods primarily target efficiency (i.e., average set size), they do not explicitly address conditional validity. In contrast, our approach is specifically designed to fill this gap by directly targeting conditional validity.
@@ 2 Methodology
In this section, we introduce our data-driven framework for learning and predicting the reliability of any given prediction interval method. Figure 1 illustrates the entire pipeline. Details are provided in the following subsections.
@@ 2.1 Asymptotic Behavior of the Conditional Coverage Probability
Figure 2 visualizes the asymptotic conditional coverage function $\eta(x)$ under several representative forms of model misspecification, illustrating that uniform convergence of $\eta_{n}(x)$ does not generally imply convergence to the nominal target level $1-\alpha$.
@@ 2.2 Estimating the Conditional Coverage Probability Function
Any probabilistic classifier capable of capturing nonlinear dependencies, such as random forests, gradient boosted trees (XGBoost), or calibrated neural networks, can be employed as the base learner for $\hat{\eta}_{n_{\text{eval}}}$. The detailed training procedure is presented in Algorithm 1 in Appendix B, and the theoretical background on calibration is provided in Appendix C.
@@ 2.3.1 The Conditional Validity Index (CVI) and its Decomposition
We first define the primary scalar metric, the Conditional Validity Index (CVI), as the mean absolute deviation of the estimated reliability from the nominal level $1-\alpha$:
@@
A lower CVI indicates that the predictor’s local coverage probabilities are tightly concentrated around the target level. Crucially, this total deviation can be decomposed into two distinct components: Undercoverage Risk ($\text{CVI}_{\text{U}}$) and Overcoverage Cost ($\text{CVI}_{\text{O}}$), such that $\text{CVI}=\text{CVI}_{\text{U}}+\text{CVI}_{\text{O}}$ as described below. Fix a tolerance parameter $\gamma\in[0,1]$.
1. Safety Assessment (Undercoverage).  These metrics quantify the risk of the model failing to meet the target coverage, which is the primary concern in high-stakes applications.
  • Undercoverage Risk ($\text{CVI}_{\text{U}}$): The average reliability shortfall as the primary safety metric.
@@
  • Undercoverage Rate ($\pi_{-}$): The proportion of samples where the estimated coverage is strictly below the target, measuring the breadth of the violation.
@@
  • Conditional Mean Undercoverage (CMU): The average shortfall calculated only among the undercovered samples. This measures the severity of the violation.
@@
2. Efficiency Assessment (Overcoverage).  These metrics quantify the cost of the model being overly conservative, which implies unnecessarily wide intervals.
@@
  • Overcoverage Rate ($\pi_{+}$): The proportion of samples where the estimated coverage is strictly above the target (breadth).
@@
  • Conditional Mean Overcoverage (CMO): The average excess coverage calculated only among the over-covered samples. This measures the degree of conservatism (severity).
@@
The tolerance parameter $\gamma$ is chosen to be a small constant close to zero and is introduced for practicality and interpretability. First, since conformal prediction methods are explicitly calibrated to target the coverage level $1-\alpha$, the estimated conditional coverage $\hat{\eta}_{n_{\textup{eval}}}(X)$ may be highly concentrated in a neighborhood of $1-\alpha$; without a tolerance margin, the ratio-based metrics $\pi_{-}$ and $\pi_{+}$ would be overly sensitive to minor estimation noise and numerical fluctuations. Second, even for an ideally calibrated method whose conditional coverage is uniformly centered at the target level, random estimation variability would lead to $\pi_{-}\approx\pi_{+}\approx 1/2$ when the exact threshold $1-\alpha$ is used, yielding undercoverage or overcoverage rate values that are misleading. Introducing $\gamma$ filters out such negligible deviations and ensures $\pi_{-}$ and $\pi_{+}$ capture substantively meaningful departures from the target coverage.
@@ 2.3.2 Visualizing Reliability: The CVP Curve
Complementing the scalar metrics, we introduce the Conditional Validity Profile (CVP) curve to provide a comprehensive visual diagnosis. Unlike summary statistics that compress information, this visualization characterizes the entire distribution of the estimated conditional coverage probabilities, revealing structural patterns of reliability.
@@
Geometric Interpretation. As illustrated in Figure 3, the geometry of the CVP curve directly corresponds to our numerical indices: With the tolerance parameter $\gamma=0$,
@@
Ideally, a perfectly conditionally valid model yields a flat CVP curve aligned with the horizontal line $y=1-\alpha$. Deviations from this reference line serve as visual diagnostics for reliability failures; for instance, a sharp drop at the left tail indicates a specific subpopulation suffering from severe undercoverage.
Figure 3: Conditional Validity Profile (CVP) Curve. This curve depicts conditional coverage for a standard conformal predictor (target $1-\alpha=0.8$). The red area reflects undercoverage risk ($\text{CVI}_{\text{U}}$), the blue area represents overcoverage cost ($\text{CVI}_{\text{O}}$), and their intersection marks the fraction of undercovered samples.
@@ 2.4 Reliability-Driven Model Selection and Deployment
which serves as an estimate of the conditional coverage at $X_{\mathrm{new}}$. In applications, $\hat{c}(X_{\mathrm{new}})$ can be compared to the nominal level $1-\alpha$ to identify inputs for which the conformal prediction set may exhibit poor conditional coverage. A complete description of the two-stage procedure is provided in Algorithm 2 in Appendix B.
@@ 4 Simulation Studies
We conduct a comprehensive simulation study to evaluate the CPA framework’s efficacy across diverse distributional scenarios, with three primary objectives: (1) to verify that estimated conditional validity recovers ground-truth performance; (2) to assess robustness to reliability estimator specification; and (3) to provide empirical deployment guidance. To this end, we first validate ranking fidelity, distributional recovery, and robustness across four synthetic scenarios (Sections 4.24.3). We then examine operational prerequisites—data allocation and sample complexity—to derive real-world guidelines (supported by Appendices IJ and synthesized in Section 4.4).
@@ 4.1.1 Data Generating Processes and Evaluation Protocol
Our evaluation adopts a “Selection-Deployment” protocol designed to emulate the lifecycle of real-world machine learning systems. For each experimental replication, we generate a selection dataset $\mathcal{D}_{\text{select}}$ ($n=2000$) and an independent held-out test set $\mathcal{D}_{\text{test}}$ ($n_{\text{test}}=2000$) which remains inaccessible during the selection phase. The selection set is randomly partitioned into a training set $\mathcal{D}_{\text{train}}$ ($50\%$) for developing candidate models and an evaluation set $\mathcal{D}_{\text{eval}}$ ($50\%$) for CPA-based reliability auditing.
To establish the ground truth for model selection, we implement a full-data refitting protocol. Specifically, while CPA performs selection using only the split data ($\mathcal{D}_{\text{train}},\mathcal{D}_{\text{eval}}$), the Oracle ranking is determined by retraining all candidate algorithms on the full $\mathcal{D}_{\text{select}}$ and evaluating their exact conditional coverage on $\mathcal{D}_{\text{test}}$. This rigorous comparison allows us to verify whether the ranking inferred from partial data successfully generalizes to the performance of the fully deployed models. Crucially, across all experimental settings, the data generating processes follow a unified location-scale structure $Y=\mu(X)+\sigma(X)\epsilon$, where $\epsilon$ is a standardized random variable independent of the covariates $X$. Building upon the established benchmarks from Lei et al. (2018), we construct four synthetic scenarios specifically designed to probe distinct conditional coverage challenges, ranging from nonlinear misspecification to severe feature dependence. These settings are summarized in Table 1, with full mathematical specifications detailed in Appendix D.
@@ 4.1.2 Benchmarked Algorithms
To evaluate the discriminative power of CPA, we benchmark nine algorithms for constructing prediction intervals ($1-\alpha=0.9$), selected to represent varying degrees of adaptivity. We include three classical baselines: OLS, Adaptive Residual Bootstrap, and Quantile Regression Forests (QRF). The core evaluation focuses on Conformal Prediction (CP) estimators, spanning from methods guaranteeing marginal validity (CP-Residual, CV+) to those designed for conditional coverage (CP-Studentized, CQR, LCP, RLCP). Implementation details and hyperparameter configurations are provided in Appendix E.
@@ 4.1.3 Oracle Ground Truth and Ranking
To benchmark the estimation fidelity of CPA, we compute the exact conditional coverage probability for each deployed predictor. Unlike empirical evaluation on finite test sets, our simulation environment allows us to access the true parameters of the data generating process (DGP). Specifically, leveraging the location-scale structure $Y=\mu(X)+\sigma(X)\epsilon$, we utilize the true conditional mean $\mu(x)$ and noise scale $\sigma(x)$.
Let $\hat{C}_{m}^{\text{full}}(x)=[\hat{l}_{m}(x),\hat{u}_{m}(x)]$ denote the prediction interval constructed by algorithm $\mathcal{A}_{m}$ after retraining on the full selection set (as described in Section 4.1). We define the Oracle Conditional Coverage for algorithm $m$, denoted as $\eta_{m}^{\text{oracle}}(x)$, as the analytical probability that the target falls within this interval using $F_{\epsilon}$, the cumulative distribution function (CDF) of the standardized noise $\epsilon$:
@@
This formulation eliminates finite-sample auditing variance, providing a deterministic measure of reliability. Based on this ground-truth probability, we quantify the performance of each algorithm $m$ using the Oracle Conditional Validity Index ($\text{CVI}_{m}^{\text{oracle}}$), defined as the mean absolute deviation on the independent test set $\mathcal{D}_{\text{test}}$:
@@ 4.2 Predictive Fidelity of Model Selection
The utility of CPA hinges on its ability to minimize the selection-deployment gap. To quantify this, we assess the rank concordance between the CPA-estimated ranking $\hat{\pi}$ (based on estimated CVI) and the ground-truth Oracle ranking $\pi^{\text{oracle}}$. We employ three complementary metrics: Weighted Kendall’s $\tau_{w}$ (global correlation), Hit@k (top-$k$ precision), and NDCG (selection utility).

### 23 Apr 2026 v2

@@ 1.2 Our Contributions
    A rigorous, flexible, and informative framework for auditing conditional coverage. CPA estimates the conditional coverage function $\hat{\eta}(x)$ using a data-adaptive probabilistic learner, providing a flexible alternative to coarse binning-based assessments and avoiding manual stratification. Beyond a single summary statistic, we decompose conditional coverage assessment into Safety (undercoverage risk) and Efficiency (overcoverage cost), yielding informative diagnostics of model failure modes. We further introduce the Conditional Validity Profile (CVP) curve, a visualization tool that summarizes the distribution of local conditional coverage probabilities.
@@
    Reliability-driven diagnosis and model selection (CC-Select). CPA provides a principled basis for diagnosing where a conformal predictor fails, thereby indicating which methodological refinements are worth exploring. It also enables the comparison of practically relevant conformal prediction methods or variants based on conditional validity rather than predictive efficiency alone; we refer to this selection framework as CC-Select. We demonstrate through extensive experiments that CC-Select consistently identifies models with superior conditional coverage performance.
@@ 1.3 Related Work
Conditional Coverage in Conformal Prediction.  Recent work has explored multiple approaches to improving or approximating conditional coverage in conformal prediction. One prominent line of work, initiated by Romano et al. (2019) with Conformalized Quantile Regression (CQR), directly models the conditional quantiles to produce adaptive intervals effective against heteroscedasticity. A second, broader strategy leverages an estimate of the full conditional distribution $P_{Y|X}$. This includes early work based on density estimates (Lei et al., 2013), as well as more flexible modern approaches that reframe the problem via discretization of the output space (Sesia and Romano, 2021) or by using localized approaches that up-weight calibration points near the test point (Guan, 2023). Recognizing the theoretical limitations of localized methods, Hore and Barber (2025) proposed RLCP, which achieves exact marginal coverage and, in a randomized sense, provides conditional coverage guarantees. Most recently, Gibbs et al. (2025) bridged the gap between marginal and conditional validity via a framework that interpolates between the two. Specifically, they employ augmented quantile regression on conformity scores to yield finite-sample coverage guarantees over a user-specified class of covariate shifts or subgroups. While existing methods offer theoretical improvements, their empirical validation is often limited to coarse, low-dimensional binning tests. Such approaches are inherently unscalable and may fail to detect subtle multivariate deviations. This highlights a crucial methodological gap: the lack of a scalable, quantitative, and objective framework for evaluating and comparing the conditional coverage properties of different methods. Our work is designed to directly address this issue.
Model Selection in Conformal Prediction.  Model selection in conformal prediction has begun to receive systematic attention only recently. Yang and Kuchibhotla (2025) explored the strategies of selection and aggregation and were the first to propose valid algorithmic procedures that provably maintain marginal coverage. Liang et al. (2024) built on Yang and Kuchibhotla (2025) by introducing ModSel-CP and ModSel-CP-LOO methods that correct for model selection bias in conformal prediction without requiring additional data splitting, ensuring finite-sample validity through full conformal frameworks. While these methods primarily target efficiency (i.e., average set size), they do not explicitly address conditional validity. In contrast, our approach is specifically designed to fill this gap by directly targeting conditional validity.
@@ 2 Methodology
In this section, we introduce our data-driven framework for learning and predicting the reliability of any given prediction interval method. Figure 1 illustrates the entire pipeline.
@@ 2.1 Asymptotic Behavior of the Conditional Coverage Probability
The above result characterizes the population target underlying CPA, but it does not imply that this target $\eta_{n}$ is necessarily easy to estimate. In general, learning the conditional coverage function for an arbitrary conformal prediction method may be as challenging as learning the full conditional law of $Y\mid X$. Fortunately, for relatively stable conformal predictors, the conditional coverage probability function may be reasonably well behaved, as we show both theoretically and numerically in Appendix B of the supplement.
@@ 2.2 Estimating the Conditional Coverage Probability Function
Any probabilistic classifier capable of capturing nonlinear dependencies, such as random forests, gradient boosted trees (XGBoost), or calibrated neural networks, can be employed as the base learner for $\hat{\eta}_{n_{\text{eval}}}$. The detailed training procedure is presented in Algorithm 1 in Appendix C, and the theoretical background on calibration is provided in Appendix D.
@@ 2.3.1 The Conditional Validity Index (CVI) and its Decomposition
We define the primary scalar metric, the Conditional Validity Index (CVI), as the mean absolute deviation of the estimated reliability from the nominal level $1-\alpha$:
@@
A lower CVI indicates that the predictor’s local coverage probabilities are concentrated around the target level. We decompose this total deviation into Undercoverage Risk ($\text{CVI}_{\text{U}}$) and Overcoverage Cost ($\text{CVI}_{\text{O}}$), so that $\text{CVI}=\text{CVI}_{\text{U}}+\text{CVI}_{\text{O}}$. Fix a tolerance parameter $\gamma\in[0,1]$.
1. Safety Assessment (Undercoverage).  These metrics quantify failure to meet the target coverage.
  • Undercoverage Risk ($\text{CVI}_{\text{U}}$): The average reliability shortfall.
@@
  • Undercoverage Rate ($\pi_{-}$): The proportion of samples whose estimated coverage falls below the target.
@@
  • Conditional Mean Undercoverage (CMU): The average shortfall among undercovered samples.
@@
2. Efficiency Assessment (Overcoverage).  These metrics quantify unnecessary conservatism.
@@
  • Overcoverage Rate ($\pi_{+}$): The proportion of samples whose estimated coverage exceeds the target.
@@
  • Conditional Mean Overcoverage (CMO): The average excess coverage among overcovered samples.
@@
The tolerance parameter $\gamma$ is taken to be a small constant close to zero. It prevents $\pi_{-}$ and $\pi_{+}$ from being dominated by negligible numerical fluctuations around $1-\alpha$, so that these rates reflect substantively meaningful deviations from the target coverage.
@@ 2.3.2 Visualizing Reliability: The CVP Curve
Complementing the scalar metrics, we introduce the Conditional Validity Profile (CVP) curve, which visualizes the distribution of estimated conditional coverage probabilities.
@@
Geometric Interpretation. As illustrated in Figure 2, the geometry of the CVP curve directly corresponds to our numerical indices. With $\gamma=0$,
@@
Ideally, a perfectly conditionally valid model yields a flat CVP curve aligned with the horizontal line $y=1-\alpha$. Deviations from this reference line diagnose reliability failures; for instance, a sharp drop at the left tail indicates a subpopulation with severe undercoverage.
Figure 2: Conditional Validity Profile (CVP) Curve. This curve depicts conditional coverage for a standard conformal predictor. The red area reflects undercoverage risk ($\text{CVI}_{\text{U}}$), the blue area represents overcoverage cost ($\text{CVI}_{\text{O}}$), and their intersection marks the fraction of undercovered samples ($\pi_{-}$).
@@ 2.4 Reliability-Driven Model Selection and Deployment
which serves as an estimate of the conditional coverage at $X_{\mathrm{new}}$. In applications, $\hat{c}(X_{\mathrm{new}})$ can be compared to the nominal level $1-\alpha$ to identify inputs for which the conformal prediction set may exhibit poor conditional coverage. A complete description of the two-stage procedure is provided in Algorithm 2 in Appendix C.
@@ 4 Simulation Studies
We conduct a comprehensive simulation study to evaluate CPA across diverse distributional regimes. The study has three objectives: (1) verify that the estimated conditional validity recovers ground-truth performance; (2) assess robustness to the choice of reliability estimator; and (3) extract practical guidance for deployment. We first study ranking fidelity, distributional recovery, and robustness across four synthetic scenarios (Sections 4.24.3). We then investigate operational requirements—data allocation and sample complexity—to derive actionable recommendations, supported by Appendices JK.
@@ 4.1.1 Data Generating Processes and Evaluation Protocol
Our evaluation adopts a “Selection–Deployment” protocol that mirrors the lifecycle of real-world machine learning systems. In each replication, we generate a selection dataset $\mathcal{D}_{\text{select}}$ ($n=2000$) and an independent held-out test set $\mathcal{D}_{\text{test}}$ ($n_{\text{test}}=2000$), which remains inaccessible during model selection. We randomly split $\mathcal{D}_{\text{select}}$ into a training set $\mathcal{D}_{\text{train}}$ ($50\%$) for fitting candidate models and an evaluation set $\mathcal{D}_{\text{eval}}$ ($50\%$) for CPA-based reliability auditing.
To establish the ground truth for model selection, we use a full-data refitting protocol. CPA performs selection using only the split data ($\mathcal{D}_{\text{train}},\mathcal{D}_{\text{eval}}$), whereas the Oracle ranking is obtained by retraining all candidate algorithms on the full $\mathcal{D}_{\text{select}}$ and evaluating their exact conditional coverage on $\mathcal{D}_{\text{test}}$. This comparison directly tests whether rankings inferred from partial data transfer to the fully deployed models.
Across all experimental settings, the data-generating processes follow the common location-scale form $Y=\mu(X)+\sigma(X)\epsilon$, where $\epsilon$ is a standardized random variable independent of the covariates $X$. Building on the benchmark constructions of Lei et al. (2018), we design four synthetic scenarios that target distinct conditional coverage challenges, from nonlinear misspecification to strong feature dependence. Table 1 summarizes these settings, and Appendix E provides their full mathematical specifications.
@@ 4.1.2 Benchmarked Algorithms
To evaluate the discriminative power of CPA, we benchmark nine prediction-interval algorithms at coverage level $1-\alpha=0.9$, chosen to span a broad spectrum of adaptivity. We include three classical baselines—OLS, Adaptive Residual Bootstrap, and Quantile Regression Forests (QRF)—and six conformal prediction (CP) methods. The CP methods range from procedures that primarily guarantee marginal validity (CP-Residual, CV+) to approaches designed to better adapt to conditional coverage variation (CP-Studentized, CQR, LCP, RLCP). Implementation details and hyperparameter configurations are provided in Appendix F.
@@ 4.1.3 Oracle Ground Truth and Ranking
To benchmark the estimation fidelity of CPA, we compute the exact conditional coverage probability for each deployed predictor. Unlike finite-sample empirical evaluation, the simulation environment gives us access to the true parameters of the data-generating process (DGP). In particular, leveraging the location-scale structure $Y=\mu(X)+\sigma(X)\epsilon$, we use the true conditional mean $\mu(x)$ and noise scale $\sigma(x)$.
Let $\hat{C}_{m}^{\text{full}}(x)=[\hat{l}_{m}(x),\hat{u}_{m}(x)]$ denote the prediction interval constructed by algorithm $\mathcal{A}_{m}$ after retraining on the full selection set, as described in Section 4.1. We define the Oracle Conditional Coverage for algorithm $m$, denoted by $\eta_{m}^{\text{oracle}}(x)$, as the analytical probability that the target falls within this interval under $F_{\epsilon}$, the cumulative distribution function (CDF) of the standardized noise $\epsilon$:
@@
This formulation removes finite-sample auditing variance and therefore yields a deterministic measure of reliability. Based on this ground-truth quantity, we summarize the performance of algorithm $m$ using the Oracle Conditional Validity Index ($\text{CVI}_{m}^{\text{oracle}}$), defined as the mean absolute deviation on the independent test set $\mathcal{D}_{\text{test}}$:
@@ 4.2 Predictive Fidelity of Model Selection
The utility of CPA hinges on whether it narrows the selection–deployment gap. To quantify this, we measure the rank concordance between the CPA-estimated ranking $\hat{\pi}$, based on estimated CVI, and the ground-truth Oracle ranking $\pi^{\text{oracle}}$. We report three complementary metrics: Weighted Kendall’s $\tau_{w}$ for global correlation, Hit@k for top-$k$ precision, and NDCG for selection utility.